OpenStudy (anonymous):

Evaluate the expression - 4a - 5b - 9 for a = 3 b = -2

pooja195 (pooja195):

- 4a - 5b - 9 \[\huge~\rm~\bf- 4(3) - 5(-2) - 9 \]

pooja195 (pooja195):

-4*3=?

OpenStudy (anonymous):

-31?

pooja195 (pooja195):

-4 times 3=?

OpenStudy (anonymous):

-12

pooja195 (pooja195):

Good now we have \[\huge~\rm~\bf~−12−(5)(−2)−9\] -5 times -2=?

OpenStudy (anonymous):

10

pooja195 (pooja195):

Good \[\huge~\rm~\bf~−12−(−10)−9 \] When we have 2 negative signs together they become a postitive \[\huge~\rm~\bf~−12+10−9 \] -12+10=?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

Good −2−9=?

OpenStudy (anonymous):

sorry -11

pooja195 (pooja195):

Perfect :)

OpenStudy (anonymous):

Thank you pooja195